OpenStudy (anonymous):

What is the least common multiple of 6 and 3? A. 3 B. 9 C. 18 D. 6

OpenStudy (anonymous):

d. 6

OpenStudy (imstuck):

6. It can't be a number less than 6 because 6 only goes into numbers equal to or greater than itself.

OpenStudy (anonymous):

thanks @IMStuck

OpenStudy (imstuck):

to do the oatmeal cookie problem, you have 3/4 of the dough with raisin, right? If you add nuts to 2/5 of the 3/4 what has nuts and raisins? The way to do this is to translate your words to an equation. The word "of" means to multiply. So 2/5 of 3/4 is\[\frac{ 2 }{ 5 }\times \frac{ 3 }{ 4 }=\]

OpenStudy (anonymous):

i think B

OpenStudy (imstuck):

What does it equal? Multiply straight across the top and straight across the bottom.

OpenStudy (imstuck):

No it's not B.

OpenStudy (imstuck):

2 times 3 is 6...5 times 4 is 20, so you have 6/20. Reduce that and what do you get?

OpenStudy (anonymous):

a

OpenStudy (imstuck):

No it's not A either.

OpenStudy (anonymous):

c

OpenStudy (imstuck):